OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = 4(1/2)^x + 1 − 3?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Y-intercept of (0, −1), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, −1), starts down on the left, gets closer to y = −3 on the right Y-intercept of (0, 1), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, 1), starts down on the left, gets closer to y = −3 on the right

OpenStudy (phi):

is this \[ y = \left(\frac{1}{2}\right)^{x+1} -3 \]?

OpenStudy (anonymous):

you forgot the 4 in front of the 1/2

OpenStudy (anonymous):

other than that, you were right

OpenStudy (phi):

\[ y = 4\left(\frac{1}{2}\right)^{x+1} -3 \]

OpenStudy (anonymous):

yes! :)

OpenStudy (phi):

the y-intercept is the y value when x is 0 any idea what y is when x=0?

OpenStudy (anonymous):

1?

OpenStudy (phi):

you erase the x and put 0 in its place in the formula what do you get ?

OpenStudy (anonymous):

ok hold on:)

OpenStudy (anonymous):

0+1=1 right?

OpenStudy (anonymous):

-1! sorry

OpenStudy (phi):

the exponent x+1 becomes 0+1 or just 1 now do (1/2)^ 1 which is just 1/2 then do 1/2 * 4 - 3 or 2 - 3 or -1

OpenStudy (phi):

follow?

OpenStudy (anonymous):

yes okay so now we can elimate some answer choices and yes i do

OpenStudy (phi):

yes, it's A or B

OpenStudy (phi):

what do you get for x= -2 any idea?

OpenStudy (anonymous):

hold on:)

OpenStudy (anonymous):

-5?

OpenStudy (phi):

x+1 becomes -2 + 1 = -1 now do (1/2)^(-1) the negative exponent means you can "flip" the fraction and make the exponent positive so you get 2^1 or just 2 now do 4*2-3 = 8-3 = 5

OpenStudy (anonymous):

Oh! I get it :)

OpenStudy (phi):

at x=2, y=5 at x=0, y= -1 starts high on the left and gets smaller (asymptotes to -3)

OpenStudy (phi):

**x=-2 , y=5

OpenStudy (anonymous):

I see, I see

OpenStudy (anonymous):

so would it be B?

OpenStudy (phi):

yes, I think B. The language is not very clear. (whoever wrote it flunked basic English)

OpenStudy (anonymous):

hahah! I totally agree<3 :D